Find the range of f(x)=-x+4 for the domain {-3,-2,-1,1}
Answer by nerdybill(7384)

Find the range of f(x)=-x+4 for the domain {-3,-2,-1,1}
.
simply substitute each of the given values of the domain for x.
-3:
f(-3)=-(-3)+4=3+4=7
.
-2:
f(-2)=-(-2)+4=2+4=6
.
-1:
f(-1)=-(-1)+4=1+4=5
.
1:
f(1)=-(1)+4=-1+4=3
.
Range is then {3, 5, 6, 7}